The area of a triangle is 15 square meters. The base is 1 meter less than the height. What is the height of the triangle?.
Answer by sachi(548)

let the ht=h m
then base b=h-1 m
so the area= 1/2*b*h=1/2*(h-1)*h=15 given
h^2-h-30=0
or, (h-6)(h+5)=0
or,h=6 or -5
so h=5 m ans